### v2

@@ Abstract
Multi-scale tile maps are essential for geographic information services, serving as fundamental components of surveying and cartographic workflows. Although existing image generation networks can efficiently produce map tiles from remote sensing images, they tend to emphasize texture features over geospatial characteristics and scale properties. Consequently, these methods are less effective in extracting vital geographic semantic information for cartography. They also fail to generate continuous multi-scale tile maps in large-format cartographic applications. This paper introduces C2GM, a novel framework for generating multi-scale tile maps through conditional guided diffusion and multi-scale cascade generation. Specifically, we develop a scale-driven conditional feature fusion encoder to extract object priors from remote sensing images and implement cascade reference double branch input, ensuring accurate representation of complex features. The generated low-level tiles serve as constraints for high-level map generation, enhancing visual continuity. Additionally, we incorporate map scale modality information using CLIP to simulate the relationship between map scale and cartographic generalization in tile maps. Extensive experimental evaluations show that C2GM consistently achieves state-of-the-art performance across all metrics, facilitating the rapid and effective generation of multi-scale large-format tile maps for emergency response and remote mapping applications. The source code, pre-trained models, and all the experimental results will be publicly available at https://github.com/Magician-MO/C2GM upon the paper’s acceptance of this work.
@@ 1 Introduction
Multi-scale tile maps, valued for their accessibility and readability, are integral to modern society, with extensive applications in emergency response, disaster management, and urban planning [1, 2]. The proliferation of online map services employing multi-scale tile maps, such as Google Maps, Bing Maps, and OpenStreetMap, has made it easy for individuals to access and utilize these resources. Although traditional mapping methods generally depend on field surveys, modern providers like Google and Microsoft now frequently employ remote sensing and mapping vehicles for data collection. However, these cartographic techniques still heavily rely on the manual acquisition of vector data, cartographic generalization and manual selection, hindering the delivery of real-time map services [3, 4]. In practical scenarios, such as mapping remote areas or conducting emergency rescue operations, the rapid generation of tiled maps and real-time updates of map services is particularly crucial. Advancements in remote-sensing technology now allow satellites, drones, and aircraft to capture high-resolution images rich in geographical information rapidly. This development makes it feasible to create multi-scale map tiles from real-time remote sensing data (Fig.1 (b)), offering a more efficient alternative to traditional cartographic workflows[5, 6, 7]. Such an approach significantly accelerates map production and ensures timely updates, particularly in inaccessible or hazardous areas for ground-based surveys.
@@
While previously considered a highly challenging task, recent advancements in image generation have transformed the field of cartographic [8]. These techniques now present a promising solution for synthesizing map tiles directly from remote sensing imagery [9, 10]. They reframing the problem as a generative task within a data-driven framework, aiding in the learning of cartographic knowledge from a wide range of tile maps as cartographic results [11].
Current map generation model, exemplified by GeoGAN [11], SMAPGAN [12], and CreativeGAN [13], utilize image translation [14, 15] methods to achieve map creation from remote sensing imagery. However, these approaches remain constrained to fixed-scale, single-layer map generation paradigms. Specifically, they exhibit insensitivity to hierarchical scale information inherent in tile-based maps and lack the capacity for multi-scale tile map production. Moreover, Multi-level map generate method can generate multi-scale maps using map-level [16, 17], but fixed level encoding fails to provide a detailed understanding of map scales and overlooks cartographic knowledge associated with scale modality information. The current flaws of applying generation method in multi-scale cartography from remote sensing images lay in three aspects:
1) Single-scale independent models are insufficient for generating coherent multi-scale maps required for hierarchical decision-making. Utilizing the cartographic expertise embedded in extensive tile map samples (see Fig. 1(a)) to enhance scale expression accuracy in generated maps remains a significant challenge. Current methodologies often rigidly link map levels to generated outputs, overlooking the integration of scale-specific characteristics including spatial resolution, map scale, and visual presentation. This oversight diminishes the model’s capacity to leverage cartographic expertise from multi-scale maps, resulting in inaccurate scale representation and compromising the practical utility of generated cartographic products.
@@
To address these challenges, we propose a novel framework for generating multi-scale tile maps from remote sensing imagery, termed Cascading Conditional Generative Cartography frameworks (C2GM). This framework leverages the principles of conditional diffusion models and multi-scale cascade generation to enhance the quality of generated maps. By integrating scale information and cascading references, C2GM effectively captures the intricate relationships between map scales and geographic features, ensuring coherent and visually appealing outputs. Unlike prior single-stage map generation methods, our self-cascaded map generation framework recursively produces multi-scale tile maps from a given spatial scene. This unified generation model is applicable across various map scales: during each stage of map generation, variables such as the low-level map from the previous stage, map level, scale, and spatial resolution are integrated as controllable parameters to guide the creation of higher-scale, high-level map tiles. As the stages progress, the produced tile maps exhibit diverse spatial resolutions, feature representations, and map syntheses, enabling the parallel generation of multi-scale map tiles for a given large-format spatial scene.
Additionally, we have developed a real-world remote sensing image dataset CSCMG (Cross-Scale Cascade Map Generation dataset) from Glasgow, UK, to validate the effectiveness of our proposed method and address limitations in existing datasets for multi-scale map generation. This dataset achieves extensive sample area coverage through coordinate sampling and includes continuous regional tile data from the London area to support continuous map range generation testing. The multi-scale tile samples within the dataset enable the verification of cross-scale map cascade generation. Furthermore, map scale information—such as map level, scale, and resolution associated with each remote sensing image-tile map sample pair—provides essential constraints for multi-scale map generation.
In summary, our main contributions are as follows.
1) We have developed a CLIP-based scale-modality encoding mechanism and constructed a multimodal cartographic feature fusion module based on SPADE. This approach dynamically achieves data-driven embedding of multi-scale cartographic knowledge priors, aligns cartographic generalization with geographic semantics, and effectively integrates the expertise of cartographic synthesis specialists into the generative model.
2) We eliminate edge discontinuities while maintaining semantic coherence across scales by employing low-level map tiles as spatial constraints in a cascaded generation strategy. It enables boundless map generation for uncharted regions, where conventional methods produce fragmented outputs unsuitable for emergency coordination.
@@ 2.1 Image-to-image generation methods
Image-to-image (I2I) translation has emerged as a pivotal technique for cross-domain visual representation learning, particularly in cartographic applications where remote sensing imagery requires transformation into structured map formats. This conversion process faces a fundamental challenge: the significant domain gap between raw sensor data and stylized map elements. To bridge this domain discrepancy, modern frameworks employ specialized architectures that explicitly model cross-domain mapping relationships. These architectures typically incorporate multi-scale feature extractors and domain-specific constraints to enhance output fidelity [18]. The objective of these frameworks goes beyond replicating the source imagery at a pixel level; it also involves capturing latent semantic representations and modeling complex feature distributions. Generally, generative models aim to synthesize artificial data instances that approximate the intrinsic probability distribution of authentic samples, based on data conforming to specific distributional constructs [19].
The initial generative model is derived from variational autoencoder (VAE), introduced by Kingma [20], pioneered probabilistic modeling for data generation. However, due to limitations in synthesizing high-resolution outputs and inherent instability during adversarial training, VAEs were superseded by GAN-based (Generative Adversarial Networks) architectures [9]. Current techniques ormulated using Conditional Generative Adversarial Networks (cGANs) [15], such as Pix2Pix [14] and Pix2PixHD [15], achieved supervised cross-domain mapping by incorporating input-output constraints. It effectively visualized semantic labels, reconstructed edges into objects, and synthesized colors. Notably, StarGAN [21] provided a paradigmatic solution for multi-domain translation through a unified architecture design, while CycleGAN [22] pioneered unpaired translation methodologies, greatly expanding the technological applicability in diverse cartographic scenarios.
Drawing inspiration from thermodynamics, diffusion models employ iterative noise introduction to capture and learn patterns of informational decay caused by noise. In turn, this enables the generation of images based on the learned dynamics [23]. These models have demonstrated exceptional performance in generating high-resolution images, with the Denoising Diffusion Probabilistic Model (DDPM) [24] being a prominent representative in the field. DDPM operates by reversing the diffusion process: in the forward process, noise is progressively added to the original data until it transforms into Gaussian noise. The model is trained to reverse this process step-by-step, reconstructing data from noise. Building on this foundation, Nichol et al. [25] introduced a learning variance strategy within DDPM, significantly improving the quality of synthesized images while enhancing the speed of the sampling process. To expedite sampling further, Song et al. proposed the Denoising Diffusion Implicit Model (DDIM) [26], which reformulated the diffusion process as non-Markovian. DDIM, without altering DDPM’s training methodology, significantly boosted sampling speed while maintaining nearly identical image generation quality. Furthermore, DDIM established a deterministic generation process, ensuring that the resulting images were solely influenced by the initial noise used in the reverse process.
The diffusion models are primarily unconditional generation models, producing images through random sampling without relying on external controls. On the other hand, conditional diffusion models generate images by incorporating additional control signals. Such models allow for the integration of diverse input conditions, spanning classification-to-image generation [27], image-to-image synthesis [28, 29, 30], text-to-image generation [31, 32, 33], and image editing [34, 35, 36, 37], among others. However, as DDPM’s reverse process operates directly in pixel space, predicting high-dimensional noise demands substantial computational resources and memory, posing challenges for high-resolution image generation. Two principal strategies have been devised to surmount this limitation: cascade generation [32, 27], which begins with the generation of a low-resolution image and progressively synthesizes higher-resolution images by treating the preceding stages as conditional inputs; and latent space generation [31], which compresses the image into a lower-dimensional latent space using an encoder-decoder framework and performs the denoising diffusion process within this latent space. Both approaches have proven effective in optimizing the efficiency of high-resolution image synthesis using diffusion models.
@@
2.2 Generation maps from remote sensing image
The generation of maps from remote sensing imagery has garnered significant attention in recent years, particularly with the advent of deep learning techniques. With the rapid advancement of remote sensing technology, capturing high-resolution remote sensing imagery has become increasingly effortless and efficient. This progress paves the way for adopting intelligent interpretation techniques based on deep learning, which inherently depend on high-quality data for optimal training. The primary tasks associated with the intelligent analysis of remote sensing imagery include semantic segmentation, feature object extraction, target detection, and scene classification. Remote sensing images are composed of many geographical elements, often intricate and diminutive, presenting notable challenges for their semantic segmentation. Existing methods for semantic segmentation in this domain have primarily focused on addressing these difficulties, particularly in the accurate segmentation of geographic elements within remote sensing imagery [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 5, 48]. In terms of the pixel-level classification of land use and land cover within remote sensing imagery, significant attention has been devoted to challenges associated with the classification of large surface areas and heterogeneous land cover types [49, 50, 51, 52, 7, 53, 5, 54]. Similarly, research aimed at feature object extraction has explored methods to detect specific elements, such as roads, by leveraging their geometric and topological characteristics for improved precision in extraction [55, 56]. For target detection, existing approaches tailored for remote sensing imagery have primarily concentrated on overcoming the challenges posed by detecting small, densely clustered objects, irrespective of their orientation or rotation [57, 58, 59, 60, 61, 62]. Meanwhile, scene classification has increasingly integrated multimodal technologies, combining visual data from remote sensing imagery with geographical textual information. This synergy has given rise to foundational image-text models within the remote sensing field [63, 64, 65, 66, 67], establishing a robust framework for a wide range of downstream applications. However, the task of generating high-quality maps from remote sensing imagery remains a complex challenge due to the inherent differences between the two domains.
Some researchers have using generate techniques to generate maps from remote sensing imagery: facilitate transitions from remote sensing imagery to map tiles through the adversarial training of generators and discriminators [14]. However, these approaches predominantly emphasize the intrinsic features of images while overlooking the relevance and complementarity of geographical features within remote sensing images, resulting in maps that inadequately represent complex geographical features. Various GAN-based map generation methods have improved map generation to some extent. For instance, GeoGAN [11] employs a Conditional Generative Adversarial Network (cGAN) [15] to translate remote sensing imagery into map representations, while SMAPGAN [12] leverages incompletely paired remote sensing and internet map sample data to produce maps using semi-supervised generative, although it does not fully overcome the issue of map quality. Additionally, CreativeGAN [13] incorporates semantic information to improve the accuracy of generated maps. These GAN-based methodologies refine the outputs through adversarial training dynamics between generators and discriminators, where discriminators assess image authenticity and penalize deviations from the target appearance. However, these methods predominantly focus on single-level map generation and fail to accommodate the intricate requirements of multi-scale map generation.
In emergency mapping, there is a necessity for small-scale maps to provide an overview with general geographic information, alongside large-scale maps to precisely depict local geographic changes. To tackle the specific challenges associated with multi-scale map generation, Chen et al. [68] proposed a straightforward training approach for producing hierarchical maps. At the highest level ($k$-th level, where $k$ denotes the maximum scale), this strategy generates maps directly from remote sensing imagery, whereas at lower levels, the mappings between the $k$-th and $(k-1)$-th levels are learned through independently trained models. However, this approach exhibits notable limitations: the training process is complex, as each level requires a distinct model, and the accuracy of the highest-level map constrains the quality of lower-level maps. This results in insufficient utilization of the features within remote sensing imagery. In contrast, CscGAN [16] developed a multi-scale map generation network capable of producing multi-level maps using a single model. This approach introduces a map-level classifier, inspired by StarGAN [21], as an extension to the discriminator. It utilizes preprocessed road morphology data as input. Nevertheless, this method struggles to generate consistent multi-scale maps solely from remote sensing imagery, as its multi-level maps pertain to disparate regions rather than representing varying scales of the same area. This limitation narrows its applicability for multi-scale map generation within a cohesive geographic region.
Building upon these advancements, Fu et al. [17] currently represent the state-of-the-art through their level-aware generation framework. They proposed a multi-level map generation network that employs a level classifier to distinguish and generate map tiles across different levels. Their network introduces a map element extractor to capture significant geographic features from satellite imagery and incorporates a multi-level fusion generator to produce seamless multi-level maps from initial lower-level map inputs. However, their architecture relies on independently generated map tiles, which, when assembled, often exhibit visual discontinuities that detract from the overall readability and aesthetic coherence of the resulting map.
In conclusion, existing methodologies for map generation often focus solely on single-level maps or fail to meet the requirements for consistent multi-scale map production. These approaches frequently overlook the complex relationships between map scales and geographic features, resulting in suboptimal outcomes. Furthermore, many methods rely on independent generation processes for each map tile, which leads to visual discontinuities and a lack of coherence across the generated maps. These limitations impede the practical application of such methods in real-world scenarios, particularly in emergency mapping and large-format cartographic applications. To address these limitations, an end-to-end generative approach for cartography has been proposed, specifically designed to produce high-quality, seamless, and wide-area multi-scale tile maps from remote sensing imagery. This method effectively tackles the challenges of multi-scale tile map generation, ensuring improved consistency and usability for regional map production.
Figure 2: The overall structure of C2GM encapsulates a hierarchical, self-cascading design aimed at producing multi-scale, large-area tile map for specified spatial scenarios. The generation process is divided into multiple stages, starting with lower-level maps and gradually transitioning to higher-level ones. In each stage, the generation process depends on the lower-level tile previously produced and its scale information. This multi-stage generation strategy not only enhances the quality of map tile generation but also ensures consistency and coherence across different tiles by learning the disparities between various map scales and integrating cartography expertise.
@@ 3.1 Overall framework
We propose a novel self-cascading generative mapping framework, designed to generate tile map across diverse scales. As depicted in Fig. 2, the overarching design of C2GM enables the recursive reconstruction of large-scale tile map using a unified generative model.
The generation process is structured into sequential stages: initially synthesizing lower-level maps, then progressively creating higher-level ones through iterative refinement. The model incorporates three key inputs at each generative stage: remote sensing images, cascading references, and corresponding scale information. Remote sensing images provide essential geo-object features for cartographic synthesis; lower-level map tiles serve as cascading references, offering pre-existing cartographic characteristics to guide the generation process; and scale information enables the model to adapt to and represent map features at various scales. Unlike auto-regressive models prioritizing resolution enhancement, our cascading framework incorporates scale-specific cartographic knowledge. This design allows the model to learn hierarchical geographic representations, ensuring that higher-level maps generalize features such as road simplification and building aggregation by cartographic standards instead of merely refining pixel-level details.
At the $k$th generation stage, let $x_{0}^{(k)}$ denote the target map and $s^{(k)}$ represent the scale information, encompassing elements such as map layer $s_{z}^{(k)}$, spatial resolution $s_{r}^{(k)}$, scale ratio $s_{s}^{(k)}$, and geospatial representation $s_{f}^{(k)}$. When generating the map for the $(k+1)$th stage, the input conditional variables $x_{0}^{(k)}$ and $s^{(k)}$ are embedded into the model. Assuming that during each stage, the model increases the tile level $s_{z}^{(k)}$ of the input tile by $n$ levels (e.g., $n=1$), and if the map tile size generated at the $k$th stage is $H\times W$ pixels, the $(k+1)$th stage output will comprise $4^{n}$ map tiles at level $s_{z}^{(k)}+n$. These new tiles correspond to the geographic area covered by the $k$th stage tile but are represented within a frame of $2^{n}\cdot H\times 2^{n}\cdot W$ pixels.
Through a parallel generation mechanism, the above operation can be iteratively repeated $m$ times, starting from the $k$th stage, eventually resulting in a full tile map consisting of multiple layers and scales. Formally, the resulting tile map can be expressed as:
$\mathcal{X}=\{x_{0}^{(k)},x_{0}^{(k+1)},\dots,x_{0}^{(k+m)}\},$
@@
where the tile map produced at the $(k+m)$th stage, $x_{0}^{(k+m)}$, corresponds to the map layer $s_{z}^{(k)}+m\cdot n$ and has an ultimate size of $H\cdot 2^{(n^{m})}\times W\cdot 2^{(n^{m})}$ pixels.
3.2 Conditional denoising in map generation
The Denoising Diffusion Probabilistic Models (DDPMs) represent a class of advanced generative models rooted in the principles of the diffusion process. The fundamental mechanism of DDPMs involves the gradual addition and subsequent removal of noise, simulating a diffusion-like transformation of data to generate new samples [24]. These two phases—forward diffusion and reverse denoising—correspond to the training and sampling stages of the model, respectively. The traditional DDPM conducts both the forward and reverse processes directly over the entire pixel space, with the diffusion process mathematically described as follows:
$y_{t}=\sqrt{\bar{\alpha_{t}}}y_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon,\ \ \ \ \epsilon\sim N\left(0,1\right)$ (1)
where $y_{0}$ denotes the data sampled from the true data distribution, and $\bar{\alpha}_{t}$ represents a predefined hyperparameter governing the noise schedule. The term $\epsilon$ corresponds to random noise drawn from the standard Gaussian distribution. The DDPM framework centers on training a denoising function $\epsilon_{\theta}$, which estimates the noise component embedded in $y_{t}$, constrained by an $L_{2}$ loss function:
$L=\mathrm{\mathbb{E}}_{\left(y_{0},\epsilon,t\right)}\parallel\epsilon-\epsilon_{\theta}\left(y_{t},t\right)\parallel^{2}$ (2)
In this equation, $t$ denotes an integer time step drawn from the interval $[0,T]$, while $\epsilon_{\theta}$ signifies the noise predictor that undergoes training. Minimizing the aforementioned loss function allows the model to progressively transform random noise into high-quality data throughout the reverse denoising procedure. For more precise and controlled data generation, supplementary conditioning variables can be integrated into the noise predictor, yielding an extended formulation such as $\epsilon_{\theta}(y_{t},t,c)$, where $c$ encapsulates the conditioning information. Based on the underlying principles of DDPM, each incremental step within the denoising operation can be mathematically formulated as follows:
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha_{t}}}}\epsilon_{\theta})+\sigma_{t}\epsilon,\ \ \ \ \epsilon\sim N\left(0,1\right)$ (3)
where $\alpha_{t}$, $\bar{\alpha}_{t}$, $\beta_{t}$, and $\sigma_{t}$ are predefined hyperparameters governing the noise schedule, and $\epsilon$ denotes additional noise injected during the reverse process. After $T-1$ iterations, this method yields a final sample conditioned on $c$.
Despite their high-quality results, DDPMs are computationally intensive and time-consuming for training and inference, making them less practical for certain applications. Addressing this limitation, Latent Diffusion Models shift the diffusion and denoising processes from pixel space to a lower-dimensional latent space learned by a pre-trained autoencoder. This paradigm drastically reduces the dimensional complexity while maintaining crucial data details. The encoding and decoding operations of the autoencoder can be defined as follows:
$z_{0}=\mathcal{E}\left(y\right)$ (4)
$y\approx\mathcal{D}\left(z_{0}\right)$ (5)
where $z_{0}$ denotes the latent representation of the target data $y$, $\mathcal{E}$ represents the encoder, and $\mathcal{D}$ is the decoder, both of which are pre-trained. As a result, the entire diffusion and denoising processes are executed in the latent space, where $z_{0}$ resides. This design achieves an optimal balance between computational efficiency and the preservation of data fidelity.

### v3

@@ Abstract
Multi-scale tile maps are essential for geographic information services, serving as fundamental outcomes of surveying and cartographic workflows. While existing image generation networks can produce map-like outputs from remote sensing imagery, their emphasis on replicating texture rather than preserving geospatial features limits cartographic validity. Current approaches face two fundamental challenges: inadequate integration of cartographic generalization principles with dynamic multi-scale generation and spatial discontinuities arising from tile-wise generation. To address these limitations, we propose a scale-aware cartographic generation framework (SCGM) that leverages conditional guided diffusion and a multi-scale cascade architecture. The framework introduces three key innovations: a scale modality encoding mechanism to formalize map generalization relationships, a scale-driven conditional encoder for robust feature fusion, and a cascade reference mechanism ensuring cross-scale visual consistency. By hierarchically constraining large-scale map synthesis with small-scale structural priors, SCGM effectively mitigates edge artifacts while maintaining geographic fidelity. Comprehensive evaluations on cartographic benchmarks confirm the framework’s ability to generate seamless multi-scale tile maps with enhanced spatial coherence and generalization-aware representation, demonstrating significant potential for emergency mapping and automated cartography applications.
@@ 1 Introduction
Multi-scale tile maps, valued for their accessibility and readability, are integral to modern society, with extensive applications in emergency response, disaster management, and urban planning (Skidmore et al., 1997; Ezequiel et al., 2014). The proliferation of online map services employing multi-scale tile maps, such as Google Maps, Bing Maps, and OpenStreetMap, has made it easy for individuals to access and utilize these resources (Stefanakis, 2017). Although traditional mapping methods generally depend on field surveys, modern map providers now frequently employ remote sensing and mapping vehicles for data collection. However, these cartographic techniques still heavily rely on the manual acquisition of vector data, cartographic generalization and manual selection, hindering the delivery of real-time map services (Haunold and Kuhn, 1993; Park and Yu, 2011). In practical scenarios, such as mapping remote areas or conducting emergency rescue operations, the rapid generation of tiled maps and real-time updates of map services is particularly crucial. Advancements in remote-sensing technology now allow satellites, drones, and aircraft to capture high-resolution images rich in geographical information rapidly (Shen et al., 2022; Liu et al., 2022; Zhao et al., 2024). This development makes it feasible to create multi-scale map tiles from real-time remote sensing data (Figure1 (b)). Recent advancements in image generation offer a more efficient alternative to traditional cartographic workflows(Goodfellow et al., 2014; Ganguli et al., 2019). Such an approach significantly accelerates map production and ensures timely updates, particularly in inaccessible or hazardous areas for ground-based surveys.
@@
Current map generation model, exemplified by MapGAN (Li et al., 2020), SMAPGAN (Chen et al., 2021), and CreativeGAN (Fu et al., 2022), utilize image translation (Isola et al., 2017; Wang et al., 2018) methods to achieve map creation from remote sensing imagery. However, these approaches remain constrained to fixed-scale, single-layer map generation paradigms. Specifically, they exhibit insensitivity to hierarchical scale information inherent in tile-based maps and lack the capacity for multi-scale tile map production. Moreover, Multi-level map generate method can generate multi-scale maps using map-level (Liu et al., 2021; Fu et al., 2023), but fixed level encoding fails to provide a detailed understanding of map scales and overlooks cartographic knowledge associated with scale modality information. The current flaws of applying generation method in multi-scale cartography from remote sensing images lay in three aspects:
1) Single-scale models fail to generate coherent multi-scale maps essential for hierarchical decision-making. Leveraging cartographic expertise from extensive tile map samples (Fig. 1(a)) to improve scale accuracy remains challenging. Current methods rigidly link map levels to outputs, neglecting the integration of scale-specific characteristics like spatial resolution, map scale, and visual features. This oversight limits models’ ability to utilize multi-scale cartographic knowledge, causing inaccurate scale representation and reduced practicality of generated maps.
@@
To address these challenges, we propose a novel framework, SCGM, for generating multi-scale tile maps from remote sensing imagery. By leveraging conditional diffusion models and a multi-scale cascade generation approach, SCGM effectively integrates scale information and cascading references to capture the intricate relationships between map scales and geographic features. Unlike previous single-stage methods, SCGM recursively generates multi-scale tile maps while incorporating controllable parameters, such as map scale, level, and resolution, to facilitate the production of larger-scale tiles. This unified framework enables the parallel generation of spatially consistent multi-scale tile maps for large-format scenes, as demonstrated on the CSCMG dataset — a comprehensive, real-world, multi-scale remote sensing mapping dataset that utilizes associated scale information.
In summary, our main contributions are as follows:
1) We have developed a scale-modality encoding mechanism to formalize map generalization relationships. This approach dynamically achieves data-driven embedding of multi-scale cartographic knowledge priors, aligns cartographic generalization with geographic semantics, and effectively integrates the expertise of cartographic synthesis specialists into the generative model.
2) We eliminate edge discontinuities while maintaining semantic coherence across scales by employing smaller scale map tiles as spatial constraints in a cascaded generation strategy. It enables boundless map generation for uncharted regions, where conventional methods produce fragmented outputs unsuitable for emergency coordination.
@@ 2.1 Image-to-image generation methods
Image-to-image (I2I) translation is crucial in cross-domain visual representation learning, particularly in converting remote sensing data into structured mapping outputs. Early methods predominantly employed generative adversarial networks (GANs) (Goodfellow et al., 2014; Mirza and Osindero, 2014), with pioneering works such as Pix2Pix (Isola et al., 2017), Pix2PixHD (Wang et al., 2018) and CycleGAN (Zhu et al., 2017) establishing both supervised and unsupervised frameworks for cross-domain mapping. Subsequent innovations, including StarGAN (Choi et al., 2018), expanded GAN-based architectures to accommodate multi-domain translation tasks through unified network designs. Although these approaches achieved pixel-level consistency and semantic preservation, challenges continued to arise in modeling high-dimensional feature distributions and synthesizing photorealistic details (Pang et al., 2022).
Recent diffusion-based generators mitigate key limitations of GANs — specifically, mode collapse and adversarial instability — by employing likelihood-based training, which enhances sample diversity and convergence (Sohl-Dickstein et al., 2015). DDPMs (Ho et al., 2020) pioneered this paradigm through a two-phase process: gradually corrupting data with Gaussian noise (forward) and learning to reverse it (reverse). Subsequent work improved training efficiency — Nichol et al. optimized noise variance estimation (Nichol and Dhariwal, 2021), while DDIM (Song et al., 2022) reformulated diffusion as non-Markovian, enabling faster deterministic sampling without quality loss.
Conditional diffusion models extend this framework by incorporating external guidance (e.g., class labels, text, images) for controlled synthesis (Ho et al., 2022; Saharia et al., 2022a; Rombach et al., 2022). However, pixel-space diffusion remains computationally prohibitive for high-resolution outputs. Two dominant solutions address this: cascaded generation (Saharia et al., 2022b), which progressively upscales low-resolution intermediates, and latent diffusion (Rombach et al., 2022), which operates in compressed feature spaces. Both strategies balance fidelity and efficiency, enabling scalable synthesis of complex imagery.
@@
2.2 Map generation from RS image
Researchers have employed generative techniques like GANs to enable image-to-map translation by learning pixel-level mappings from remote sensing data (Isola et al., 2017). While methods like GeoGAN (Ganguli et al., 2019), SMAPGAN (Chen et al., 2021), and CreativeGAN (Fu et al., 2022) enhance feature fidelity through conditional architectures and semantic constraints, they predominantly focus on single-scale outputs, failing to capture hierarchical geographic relationships essential for multi-scale mapping.
Efforts to address multi-scale challenges remain fragmented. Chen et al. (Chen et al., 2022) proposed hierarchical generation using separate models per scale, but error propagation from coarse-to-fine levels limits accuracy. CscGAN (Liu et al., 2021) unified multi-level generation via a single model but required auxiliary road morphology inputs, restricting scalability. The SOTA method LACG (Fu et al., 2023) introduced level-aware fusion but suffers from tile-wise inconsistencies due to independent tile generation, undermining spatial coherence in composite maps.
In conclusion, existing methods either neglect cross-scale semantic dependencies or fail to ensure visual continuity across tiles, hindering their utility in emergency mapping and large-format applications. Our work addresses these challenges through a cascaded conditional diffusion framework, enabling end-to-end synthesis of seamless, geographically consistent multi-scale tile maps from unified remote sensing inputs.
@@ 3.1 Overall framework
Figure 2: Overview of the SCGM framework: a hierarchical, self-cascading pipeline that progressively synthesizes multi-scale tile maps by refining smaller-scale outputs under explicit scale conditioning.
We propose SCGM, a self-cascading generative mapping framework for multi-scale tile map synthesis, following the architecture of LDM (Rombach et al., 2022). As shown in Figure 2, SCGM employs a unified generative model to recursively construct large-scale maps from smaller-scale representations. The framework comprises three core modules: ScaleEncoder, which embeds scale-related spatial context; MFEncoder, which fuses semantic features from remote sensing imagery and cascading references; and SFAdapter aligns multi-scale features with the denoising trajectory.
The generation proceeds in stages, with each step incorporating three inputs: remote sensing images for geographic content, lower-scale maps as spatial priors, and scale information guiding generalization. Unlike resolution-focused models (Ho et al., 2022; Tian et al., 2024), SCGM emphasizes scale-specific semantics, enabling hierarchical abstraction consistent with cartographic conventions. This structured cascade supports coherent multi-scale synthesis, with further details on feature integration and scale adaptation presented later.
At generation stage $k$, let $x_{0}^{(k)}$ denote the target map and $s^{(k)}$ the associated scale information, including map layer $s_{z}^{(k)}$, spatial resolution $s_{r}^{(k)}$, scale ratio $s_{s}^{(k)}$, and geospatial encoding $s_{f}^{(k)}$. To generate the map at stage $k+1$, the model conditions on $x_{0}^{(k)}$ and $s^{(k)}$. Assuming the model advances the tile level by $n$ steps per stage (e.g., $n=1$), an input tile of size $H\times W$ pixels at level $s_{z}^{(k)}$ results in $4^{n}$ output tiles at level $s_{z}^{(k)}+n$, collectively spanning $2^{n}H\times 2^{n}W$ pixels over the same geographic region.
This process can be iterated $m$ times, enabling parallel generation of increasingly detailed maps across multiple scales. The complete multi-scale output is expressed as:
$\mathcal{X}=\{x_{0}^{(k)},x_{0}^{(k+1)},\dots,x_{0}^{(k+m)}\}$ (1)
@@
where $x_{0}^{(k+m)}$ corresponds to level $s_{z}^{(k)}+m\cdot n$ and covers an area of size $H\cdot 2^{nm}\times W\cdot 2^{nm}$ pixels.
3.2 Conditional denoising map generation